Solve the equation:
Ally inherits $65,000 and decides to invest part of it in an education account for her daughter and the rest in a 10-year CD. If the amount she puts in the education account is $12,000 more than twice the amount she puts in the CD, how much money does Ally invest in each account?

Answer by stanbon(75887)

Ally inherits $65,000 and decides to invest part of it in an education account for her daughter and the rest in a 10-year CD. If the amount she puts in the education account is $12,000 more than twice the amount she puts in the CD, how much money does Ally invest in each account?
-----------
In Ed. account: E
In CD account:: C
-----
Equations:
E + C = 65000
E - C = 12000
---
2E = 77000
E = 38500
C = 38500+12000 = 50500
-----------
Cheers,
Stan H.
-------------